OpenStudy (anonymous):

What is a 1/4 of $460?

OpenStudy (anonymous):

115

OpenStudy (agreene):

460*0.25 = 115

OpenStudy (anonymous):

What would be another 1/4 off of that?

OpenStudy (anonymous):

1/4*the value

OpenStudy (agreene):

115*0.25 = 28.75

OpenStudy (anonymous):

1/4*115 =28.75

OpenStudy (anonymous):

1/4*460 =115

OpenStudy (anonymous):

1/4 pf 115?

OpenStudy (anonymous):

of*

OpenStudy (anonymous):

1/4X115

OpenStudy (anonymous):

I got something around 26 -_-

OpenStudy (anonymous):

no its 28.75

OpenStudy (anonymous):

check again

OpenStudy (anonymous):

Did. I got what you have.

OpenStudy (anonymous):

Thanks for the help :)